OpenStudy (anonymous):

solve the equation for y: 4x-2y=6

OpenStudy (anonymous):

y=6-4x/-2

OpenStudy (anonymous):

my answer wa y=3+2x

OpenStudy (anonymous):

or y=-3+2x Not +3

OpenStudy (anonymous):

i did't simplify it in simple term y=-3+2x .

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

you can fan me as i you?